OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 1 points) Choose the correct simplification of the expression d3 • d5. d d15 d2 d8 Question 3 (Multiple Choice Worth 1 points) Choose the correct simplification of the expression (a2)3. a5 a8 a a6 Question 4 (Multiple Choice Worth 1 points) Choose the correct simplification of the expression (8x2)3. 24x6 512x6 512x5 24x5

OpenStudy (anonymous):

please help someone i know the answers but i want to be fully sure

OpenStudy (unklerhaukus):

\[d^3 \times d^5\\=ddd\times ddddd\\=dddddddd\\=d^?\]

OpenStudy (anonymous):

d^8

OpenStudy (unklerhaukus):

\[\checkmark\]

OpenStudy (unklerhaukus):

\[(a^2)^3\\=(aa)(aa)(aa)\\=aaaaaa\\=a^?\]

OpenStudy (anonymous):

a^6

OpenStudy (unklerhaukus):

\[\checkmark\]

OpenStudy (mathstudent55):

@UnkleRhaukus Perhaps it's: (8x^2)^3

OpenStudy (unklerhaukus):

\[(8\times2)^3\\=8^3\times2^3\\=\]

OpenStudy (unklerhaukus):

hmmm, i dunno

OpenStudy (anonymous):

4,096 is the answer i keep getting

OpenStudy (anonymous):

Yes mathstudent its written the way you put it

OpenStudy (unklerhaukus):

your right @mathstudent55 that is an \(x\) and not a multiplication symbol \(\times\)

OpenStudy (mathstudent55):

\((8x^2)^3\) or \((8 \times\ 2)^3\) ?

OpenStudy (anonymous):

The first one

OpenStudy (unklerhaukus):

\[ (8x^2)^3\\=(8xx)(8xx)(8xx)\\=8\times8\times8\times xxxxxx\]

OpenStudy (anonymous):

512^6

OpenStudy (unklerhaukus):

where'd x go?

OpenStudy (anonymous):

Oh im sorry

OpenStudy (anonymous):

512x^6

OpenStudy (anonymous):

I have a couple more questions that i need help with if thats alright with you two

OpenStudy (unklerhaukus):

\[\boxed{\huge\color{red}\checkmark}\]

OpenStudy (anonymous):

ompare and Contrast: Below are two expressions. Simplify each and choose the statement that is true about each. Expression #1 Expression #2 (4x)2(x) (6x2)4 The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information. Question 6 (Multiple Choice Worth 1 points) Choose the correct simplification of the expression (c2)9. c18 c11 c512 c Question 7 (Multiple Choice Worth 1 points) Choose the correct simplification of the expression b5 • b4. b b9 b20 b−1 Question 8 (Multiple Choice Worth 1 points) Choose the correct simplification of the expression (xyz2)4. x5y5z6 x5y5z8 xyz16 x4y4z8 Question 9 (Essay Worth 2 points) Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)2

OpenStudy (mathstudent55):

\( (8x^2)^3 \) \(=(8x^2) \times\ (8x^2) \times\ (8x^2) \) \(=(8xx)(8xx)(8xx)\) \(=8 \times\ 8 \times\ 8 \times\ xxxxxx \) \(=512x^6\)

OpenStudy (anonymous):

you are right

OpenStudy (unklerhaukus):

are there your expressions, or am i reading them wrong? Expression #1 \((4x)2^x\) Expression #2 \( (6x^2)^4\)

OpenStudy (anonymous):

You are right

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

do you have a option for 5 that makes sense to you?

OpenStudy (unklerhaukus):

do you know how big \(x\) is?

OpenStudy (anonymous):

I believe it may be thisThe exponents on Expression #2 are greater than the exponents of Expression #1.

OpenStudy (anonymous):

I believe the x is suppose to turn into a one

OpenStudy (unklerhaukus):

why would \(x\) only equal one?

OpenStudy (anonymous):

Because when there is not a number for x i believe it turns into a 1 undercover correct if i'm wrong because i have never been to good with math

OpenStudy (unklerhaukus):

I see no reason that \(x\) must be one. \(x\) could be any number

OpenStudy (anonymous):

Okay your right I forgot but what does this have to do with the question for number 5

OpenStudy (unklerhaukus):

since we do not know the value for \(x\) it could be anynumber.

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

Choose the correct simplification of the expression (xyz2)4. x5y5z6 x5y5z8 xyz16 x4y4z8 Question 9 (Essay Worth 2 points) Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)2

OpenStudy (unklerhaukus):

.

OpenStudy (anonymous):

Thats all i need now i got the others correct

OpenStudy (unklerhaukus):

\[(xyz^2)^4=(xyzz)(xyzz)(xyzz)(xyzz)\]

OpenStudy (unklerhaukus):

\[=xxxxyyyyzzzzzzzz\]

OpenStudy (unklerhaukus):

=?

OpenStudy (anonymous):

x^4y^4z^8

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]

OpenStudy (anonymous):

Okay now what about the last one

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)2

OpenStudy (mathstudent55):

Part 1. To raisa a product to a power, you must raise every factor of the product to the power. To raise a power to a power, multiply the powers togehter. Part 2. \((3x^3y^4)^2\) \(= (3)^2(x^3)^2(y^4)^2\) \(= 9x^6y^8\)